OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3. −21 −16 −10 −1 Can you please help me out with this question? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is another (possibly more convoluted) way of asking you to solve 3 = 4(x+3) - 5

jimthompson5910 (jim_thompson5910):

this is because when you solve for the inverse, you end up swapping x and y (and then solving for y)

OpenStudy (anonymous):

oh ok i didnt pay attention to the question.. i thought they were asking to solve for f(x) and that's it

jimthompson5910 (jim_thompson5910):

that's if they asked you to find f(3)

OpenStudy (anonymous):

yeah, i understand but i'm kind of having trouble solving for inverses

jimthompson5910 (jim_thompson5910):

do you know how to solve 3 = 4(x+3) - 5?

OpenStudy (anonymous):

yes i do

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

19

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

how? i plugged it 3 for the x which is 4(3+3) -5 so whats inside the parantheses is 6 and 4 times 6 is 24 so 24-5 = 19

jimthompson5910 (jim_thompson5910):

no you don't plug 3 into x

OpenStudy (anonymous):

oh then i was doing this wrong the whole time.. wow lol

OpenStudy (anonymous):

so can u help me get this whole thing plz?

jimthompson5910 (jim_thompson5910):

3 = 4(x+3) - 5 3 = 4x+12 - 5 3 = 4x+7 4x+7 = 3 do you see how to finish up?

OpenStudy (anonymous):

how did u get 7?

jimthompson5910 (jim_thompson5910):

12-5 = 7

OpenStudy (anonymous):

oooh oh ohh ok lol thank you so much! that was very helpful!

jimthompson5910 (jim_thompson5910):

you're welcome